OpenStudy (anonymous):

x-3y=-12 2x+3y=11 solve by using elimination

OpenStudy (anonymous):

x-3y=-12 2x+3y=11 3x 0 = -1 3x=-1 x=-1/3

OpenStudy (texaschic101):

x - 3y = -12 2x + 3y = 11 --------------add 3x = - 1 x = -1/3 now sub -1/3 in for xin any of the original equations x - 3y = -12 -1/3 - 3y = -12 (multiply the equation by 3 to get rid of the fractions) -1 - 9y = - 36 (add 1 to each side) -9y = -36 + 1 -9y = -35 y = -35/-9 y = 35/9 check... 2x + 3y = 11 2(-1/3) + 3(35/9) = 11 -2/3 + 105/9 = 11 (common denominator is 9) -6/9 + 105/9 = 11 99/9 = 11 11 = 11 (correct) x = -1/3 y = 35/9